OpenStudy (anonymous):

2 questions. What is the solution to the equation 2(m-8)=4(m+6)? m = ____? The product of 9 and a number is 12 less than three times that number. What is the number?

OpenStudy (eric_d):

expand LHS & RHS first

OpenStudy (eric_d):

then shift the unknown m frm LHS to RHS

OpenStudy (anonymous):

2(m-8)=4(m+6)? Distribute on both sides 2m -16= 4m + 24 -2m -2m -16 = 2m +24 -24 -24 -40 = 2m /2m /2m -20 = m

OpenStudy (eric_d):

Good job @Christineluna

OpenStudy (anonymous):

The product of 9 and a number is 12 less than three times that number. 9n -12 = 3n -9n -9n -12 = -6n /-6 /-6 2= n check: 9(2) -12= 3(2) 18 -12 = 6 6=6